# Particle scale anisotropy controls bulk properties in sheared granular materials

### v1

@@ Abstract
The bulk dynamics of dense granular materials arise through a combination of particle-scale and mesoscale effects. Theoretical and numerical studies have shown that collective effects are created by particle-scale anisotropic structures such as grain connectivity (fabric), force transmission, and frictional mobilization, all of which influence bulk properties like bulk friction and the stress tensor through the Stress-Force-Fabric (SFF) relationship. To date, establishing the relevance of these effects to laboratory systems has remained elusive due to the challenge of measuring both normal and frictional contact forces at the particle scale. In this study, we perform experiments on a sheared photoelastic granular system in an quasi-2D annular (Couette) cell. During these experiments, we measure particle locations, contacts, and normal and frictional forces vectors during loading. We reconstruct the angular distributions of the contact and force vectors, and extract the corresponding emergent anisotropies for each of these metrics. Finally, we show that the SFF relation quantitatively predicts the relationship between particle scale anisotropies, the stress tensor components, and the bulk friction coefficient, capturing even transient behaviors. As such, this method shows promise for application to other dense particulate systems where fabric anisotropy can provide a useful measure of bulk friction.
@@
Figure 1: (a) Top view of photoelastic particles viewed through a darkfield polariscope (white = higher stress) inside of an annular shear cell with inner boundary rotating at a constant velocity $v$ and fixed outer boundary. Each particle provides a measurement of its own contact forces, which we measure relative to the contact position vector $r$. (b) Sketch of the interparticle coordinate system, with the branch vector $l^{ab}$ pointing from the center of particle $a$ to the contact point with particle $b$. The force exerted by particle $a$ on $b$ is given by $f^{ab}$; this can decomposed into normal $f_{n}$ and tangential $f_{t}$. The angle $\theta$ defines the relative angle between $l$ and $r$ for the branch vector and position vector of the contact between particles $a$ and $b$.
@@
Because fabric strongly influences bulk measures, even when the particles undergo identical loading history, jammed systems additionally exhibit an anisotropy in shear modulus [9, 10]. In rheological contexts, bulk friction coefficients have been shown to depend on the combination of fabric and force [11]. Theoretical micromechanical models have been developed to further understand granular material from rheology [12, 13, 14] to stability and jamming [15, 16] and many studies have quantified the anisotropic structures appearing while under load  [17, 18, 19, 20, 21].
@@
All of these challenges can be addressed with a framework first pioneered by Rothenburg and Bathurst [22], the Stress-Force-Fabric (SFF) relation connects the bulk stress tensor to particle scale anisotropies in contact orientation, contact vector length, and force vectors; it has been verified extensively in DEM simulations of granular materials of varying complexity to understand the role of particle scale structure in bulk measurements [23, 24, 25, 26]. Further work by Li and Yu [27] rigorously derived the SFF relation using a directional statistics approach and established the key assumptions. However, the SFF relation has seen little attention in laboratory systems because it requires measurements of individual particle contacts and vector forces. Recent developments in photoelastic techniques [28, 29, 30], combined with high-resolution photography now allows for force and contact resolution suitable to test the SFF relation using photoelastic granular assemblies.
@@
In this study, we use photoelastic disks  [28, 29] arranged in a single layer within an annular shear cell (Fig. 1(a)) to map emergent particle scale anisotropic structures to bulk friction and the stress tensor using the SFF relation and further apply it to capture dynamic behaviour. In doing so, we address the open question of the influence of particle scale anisotropy during granular dynamics and show the first laboratory demonstration of the SFF relation. The particles begin in an unstrained state with no observable forces and the inner boundary is rotated to shear the assembly. During these experiments, we record particle locations, contacts, and normal and frictional forces vectors for each particle and calculate the stress tensor and anisotropies from the packing information.
@@
We performed experiments in a constant-area quasi-two-dimensional annular shear cell containing $N$ = 1810 bidisperse (Vishay PhotoStress PSM-4) circular disks with an equal fraction of small particles with radii $R_{s}$ = 4.5 mm and large particles with radii $R_{l}$ = 5.5 mm ($R_{l}/R_{s}=1.2$). The particles are dusted with baking powder to reduce basal friction between the particles and the supporting sheet, and the interparticle friction coefficient is estimated to be 0.3. The inner and outer boundaries of the annulus have particle-sized cavities to promote gripping between the boundaries and the packing while under shear. The inner boundary of the annulus is rotated at a constant rate of 1 rotation every 2.5 hours and the packing undergoes simple shear. Further details of the annulus can be found in previous works [31, 32]. Using the annular strain rate $\dot{\gamma}=\frac{1}{2}\left(\frac{\partial v}{\partial r}-\frac{v}{r}\right)$ as described by Tang et al. [33], and integrating over time with appropriate boundary conditions, we calculate the strain $\epsilon$. The packing is illuminated with both red unpolarized light (particle-detection) and green circularly polarized light (darkfield polariscope). Images are taken at a frame rate of 5 seconds per frame with a Nikon D850 camera. We can extract particle position, contact and vector force data from the images of the particles as outlined in previous works  [28, 29] by using the open source code, PhotoElastic Grain Solver [30] that we modified to detect particle-particle contacts with force magnitudes with a resolution of $5\times 10^{-5}$ N.
Figure 2: Normalized angular distributions of the (a) contacts $E^{c}$, (b) normal forces $f_{n}$, and (c) tangential forces $f_{t}$ at increasing levels of strain $\epsilon$ (dark purple to yellow). The contacts are normalized by the average number of contacts $E_{0}$, and the forces are normalized by the average normal force $f_{0}$ at each given strain. Sample angular histograms of each anisotropic metric are shown in the respective insets, which have been averaged over the strain range shown in the main panel. $\theta=0$ corresponds to the radial direction, and $\theta=\pi/2$ corresponds to the azimuthal direction. Fits to the respective Fourier expansions are indicated by the solid, sinusoidal line (Eqs. (4a)-(4c)). The normalized contacts and normalized normal forces show a decrease in anisotropy with increasing strain.
@@
From the particle scale information, we calculate the stress tensor in two dimensions using the micro-structural definition [34, 35]. We compute the two dimensional stress tensor via the summation of the outer product of the force $\vec{f}$ and branch vectors $\vec{l}$ [35, 27, 22],
$\sigma_{ij}=\frac{1}{S}\sum_{a\neq b}\vec{l^{ab}_{i}}\vec{f^{ab}_{j}},$ (1)
@@
where in our two dimensional system, $S$ is the area of the annulus. Here we follow notation typical to the soil mechanics community, where the branch vector points from the center of the grain to the contact point, and the force vector points from the contact point in the direction of the contact force. Figure 1(b) shows a schematic of the vectors for two sample particles. From this formulation of the stress tensor, we are able to calculate the bulk friction coefficient,
@@
where the numerator $\tau$ is the shear component of the stress tensor, and the denominator $P$ is the pressure.
To consider the anisotropies in a granular material, the SFF relation relies on decomposing the stress tensor into angular distributions according to the branch vector normal $\bm{\hat{n}}$. Assuming the thermodynamic limit, the stress tensor can be written as
$\sigma_{ij}=\frac{zN}{S}\oint_{\theta}e^{c}(\bm{\hat{n}})\left<l_{i}f_{j}\right>|_{\bm{\hat{n}}}d\theta,$ (3)
where $z$ is the average contact number, $N$ is the number of particles, $\theta$ is the angle corresponding to the normal vector relative to the position vector of the contact $r$, and $e^{c}(\bm{\hat{n}})$ is the probability distribution of a contact appearing at a given normal vector. The SFF relation then takes the contact vector probability, average branch vector, and average force vector as a function of the contact angle, and expands it through a tensorial Fourier expansion. In the given system, which is composed of round particles, there is no directionality to the branch vector and therefore the branch vector is a scalar value determined by the average distance between two particles $l_{0}=(N_{s}R_{s}+N_{l}R_{l})/N$. Li and Yu [27] found that through the Fourier expansion the first rank Fourier coefficient is much larger than subsequent coefficients and the components of the stress tensor can be approximated as
@@
$f_{n}(\theta)$ $=f_{0}\left[1+a_{n}\cos 2\left(\theta-\theta_{a}\right)\right],$ (4b)
$f_{t}(\theta)$ $=f_{0}a_{t}\sin 2\left(\theta-\theta_{a}\right),$ (4c)
@@
where $E^{c}(\theta)$ is the probability density of contacts at a given angle, $\theta_{a}$ is the angle corresponding to the principle stress direction, $f_{0}$ is the average magnitude of the normal force, and $\theta$ is the relative angle between contact position vector and the branch vector (see Figure 1(b)). $a,a_{n},a_{t}$ are the leading order Fourier modes for the contact, normal force and tangential force distributions, respectively and can be considered as amplitudes of the anisotropy for the system.
@@
Examples of the three distributions presented in Eqs 4a)-4c) are shown in Figure 2 with the contact distribution, normal forces and tangential forces in panels (a)-(c). Sample histograms averaged over strain $\epsilon=0-1$ are shown in the inset as polar histograms. We plot the change in these anisotropic measures as a function of relative angle between the radial direction and branch vector, where $\theta=0$ indicates alignment with the radial direction and $\theta=\pi/2$ indicates alignment with the azimuthal direction. We expect that the primary direction of the normal forces and contacts to align with the $\theta\sim\pi/4$ direction as the load is distributed through the shear band. This is true for the normal forces, indicating that the majority of the load is transmitted in the shear direction. The contact distribution follows a similar trend, however there is an additional secondary feature that appears as a bump aligned in the azimuthal direction $\theta=\pi/2$. Two possible reasons for the increased number of contacts near the azimuthal direction is either due to an artifact during contact detection, or an indication of a higher order Fourier mode, possibly present due to the constant area boundary condition. Regardless, this feature is small relative to the main signal, and we use only one Fourier mode to describe the data. The tangential distributions show only a weak signal above the noise, indicating that there is no strong preference for average tangential forces to be either clockwise (purple) or counterclockwise (pink) from the contact normal.
To quantify how the anisotropies evolve under ongoing strain, we consider the angular distributions at varying levels of strain. Figure 2 shows the normalized contact probability $E_{c}$, normalized normal force distribution $f_{n}$, and the normalized tangential forces $f_{t}$ as a function of contact angle. The contacts are normalized by the mean number of contacts $E_{0}$, and the forces normalized by the average normal force $f_{0}$. The strain ranges from $\epsilon=0-1$ indicated from purple (dark) to yellow (light) in increments of 0.04. As the assembly is strained, the amplitude of the sinusoidal shape decreases for the contacts (Fig. 2(a)) and normal forces (Fig. 2(b)).
Following the method of Li and Yu [27], we expand the stress tensor (Eq. 3) with the Fourier expansions of the three anisotropic metrics (Eqs. 4a-4c) to compare the particle scale anisotropies to bulk metrics. The SFF relation states that the stress tensor for a granular packing of isotropic particles in equilibrium can be written in terms of three main metrics: the angular contact angle distribution, and the average normal forces and average tangential forces:
@@
$\sigma_{ij}=\frac{zN}{2S}l_{0}f_{0}\,\Biggl{[}\left(1+aa_{n}\right)\delta_{ij}\\ +\frac{1}{2}\left(a+a_{n}+a_{t}\right)\begin{pmatrix}\cos\theta_{a}&\sin\theta_{a}\\ \sin\theta_{a}&-\cos\theta_{a}\end{pmatrix}\Biggr{]}$ (5)
and this expansion of the stress tensor may be used to calculate the bulk friction coefficient $\mu$. Inserting Eq. 5 into Eq. 2 results in a simple sum rule:
@@
which directly relates the bulk friction coefficient to the anisotropy measures, indicating that the resistance to deformation in a granular material simply depends on the amount of anisotropy that is present in the granular assembly.
Figure 3: Panels (a)-(c) display the leading term Fourier coefficients representing the angular distribution of contacts ($a$), average normal forces ($a_{n}$), and tangential forces ($a_{t}$), respectively, within a strained granular packing under shear. These coefficients quantify the degree of anisotropy observed in each metric within the packing. Panel (d) illustrates the bulk friction coefficient of the sheared granular packing as a function of shear, with the summed anisotropies overlaid in black and the standard error in the fit represented as grey filled bars spanning the best fit. The fabric anisotropy $a$ is plotted with a black dashed line for comparison. The blue panel highlights the strain $\epsilon$ depicted in panels (a)-(c). Panel (e) presents the four components of the strain tensor, $\sigma$, as measured by Eq. 1, with principal and secondary stress axes oriented radially and azimuthally, respectively. The solid lines represent the stress tensor calculated from the anisotropies using Eq. 5.
@@
To measure the level of anisotropy for the three metric values, we fit the angular distribution of contacts, normal forces and tangential forces (Fig. 2) to the predicted Fourier decomposition in Eqs. 4a-(c). We take the principle stress direction $\theta_{a}$ and the amplitudes $a,a_{n}$ and $a_{t}$ to be free fit parameters, where the anisotropy corresponds to the amplitudes. The anisotropies are plotted as a function of strain in Fig. 3(a)-(c); note that during the initial strain there is a strong decrease in value for all three metrics indicating a decrease in overall anisotropy. We also note that the relative magnitude of $a$ and $a_{n}$ are comparable which is consistent with simulated systems of disks [36, 22], and $a_{t}$ is an order of magnitude smaller. This indicates that the tangential forces, and consequently the inter-particle friction, has only a small effect in our system on the bulk friction, due to $a_{t}$ being smaller than those reported in simulations [37, 36].
@@
Comparing the two sides of Eq. 6 (Figure 3(d)), we see that there is excellent agreement between the two definitions of the stress tensor, the particle and bulk for a complete rotation of the inner boundary. For the annular shear cell used in these experiments, we expect $\mu$ to initially decrease as it undergoes strain softening before reaching a steady state [38]. Based on the SFF relation, this indicates that the fabric and forces are initially more strongly aligned before becoming more isotropic with increased strain. From the bulk definition, we expect that initially the assembly is resistant to shear before yielding after particle rearrangement. Furthermore, the anisotropies and bulk measurements also match each other well through large rearrangements and stick slip events, indicated by large spikes in $\mu$. Notably, $\mu$ is not solely determined by the number of force chains (see supplemental video [39]), as there are strong fluctuations in the number of force chains for a given value of $\mu$. We note that the SFF relation was developed for static loading conditions and although the dynamics in our systems are slow, the inertial number is small but non-zero ($I\sim 10^{-8}$) and there could be kinetic stress present in the system.
We gain further information from the simple shear packing by comparing the particle scale stress tensor with the bulk stress tensor. We plot the stress tensor components as measured by the force-moment tensor, as a function of strain, in Fig 3(e). The radial and azimuthal stress increase during straining, and then fluctuate as the particles undergo rearrangements during stick slip events. The off axis components of the stress tensor remain small. Overlaid on the calculated stress tensor components are the predictions for the stress tensor based on the anisotropic measures based on Eq. 5. These agree well with the measured values, however there tends to be a small discrepancy between the measured and predicted values. We attribute this discrepancy to rotations in the primary stress direction, especially during initial shear.
@@
To explore the applicability of the SFF relation in the absence of photoelastic data, in Fig. 3(d) we show the effect of the fabric anisotropy alone by plotting only $a$ as a function of strain (dashed black line). $a$ follows the overall pattern of $\mu$, including the initial transient and stick-slip events, but underestimates bulk friction by roughly 30%. Because $a$ and $a_{n}$ are typically of the same magnitude, while the tangential component is small, using the fabric anisotropy alone provides a good estimate of bulk friction.
@@
In this study, we have shown experimental validation of the SFF relation using photoelastic particles in an annular shear cell. Success of the SFF relation indicates that the resistance to deformation of a granular material arises due to anisotropies in fabric and forces. Importantly, the SFF relation remains valid even as transient stress builds up and releases through stick slip events, far beyond the static loading for which the framework was initially developed. We note that this framework does not address the role of mesoscale structure and therefore does not encompass prediction of rigidity [5]; however, the SFF relation is a promising pathway to address the open challenge of connecting fluid-like granular theories [13, 12, 14] to quasi-static granular behaviour. Finally, the framework may additionally provide a way to link fabric anisotropy to bulk friction in experimental systems for which vector contact forces cannot be measured, as illustrated by the success of the contact anisotropy $a$ alone. We note that in the absence of force information, the geometric anisotropy provides a good estimate of bulk friction and could be easily applied to other deformable assemblies, including cellular tissues, foams, and dense colloid suspensions.

### v2

@@ Abstract
The bulk dynamics of dense granular materials arise through a combination of particle-scale and mesoscale effects. Theoretical and numerical studies have shown that collective effects are created by particle-scale anisotropic structures such as grain connectivity (fabric), force transmission, and frictional mobilization, all of which influence bulk properties like bulk friction and the stress tensor through the Stress-Force-Fabric (SFF) relationship. To date, establishing the relevance of these effects to laboratory systems has remained elusive due to the challenge of measuring both normal and frictional contact forces at the particle scale. In this study, we perform experiments on a sheared photoelastic granular system in an quasi-2D annular cell. During these experiments, we measure particle locations, contacts, and normal and frictional forces vectors during loading. We reconstruct the angular distributions of the contact and force vectors, and extract the corresponding emergent anisotropies for each of these metrics. Finally, we show that the SFF relation quantitatively predicts the relationship between particle scale anisotropies, the stress tensor components, and the bulk friction coefficient, capturing even transient behaviors. As such, this method shows promise for application to other dense particulate systems where fabric anisotropy can provide a useful measure of bulk friction.
@@
Figure 1: (a) Top view of photoelastic particles viewed through a darkfield polariscope (white = higher stress) inside of an annular shear cell with inner boundary rotating at a constant velocity $v$ and fixed outer boundary. Each particle provides a measurement of its own contact forces, which we measure relative to the radial polar coordinate $R$. (b) Sketch of the interparticle coordinate system, with the contact position vector $r^{ab}$ pointing from the center of particle $a$ to the contact point with particle $b$. The force exerted by particle $a$ on $b$ is given by $f^{ab}$; this can decomposed into normal $f_{n}$ and tangential $f_{t}$. The angle $\theta$ defines the relative angle between $r^{ab}$ and $R$ for the branch vector and position vector of the contact between particles $a$ and $b$.
@@
Because fabric strongly influences bulk measures, even when the particles undergo identical loading history, jammed systems additionally exhibit anisotropy in the shear modulus [9, 10]. In rheological contexts, bulk friction coefficients have been shown to depend on the combination of fabric and force [11, 12]. Theoretical micromechanical models have been developed to further understand granular material from their rheology [13, 14, 15] to stability and jamming [16, 17] and many studies have quantified the anisotropic structures appearing while under load  [18, 19, 20, 21, 22].
@@
All of these challenges can be addressed with a framework first pioneered by Rothenburg and Bathurst [23]. The Stress-Force-Fabric (SFF) relation connects the bulk stress tensor to particle scale anisotropies in contact orientation, contact vector length, and force vectors; it has been verified extensively in DEM simulations of granular materials of varying complexity to understand the role of particle scale structure in bulk measurements [24, 25, 26, 27, 28]. Further work by Li and Yu [29] rigorously derived the SFF relation using a directional statistics approach and established its key assumptions. However, the SFF relation has seen little attention in laboratory systems because it requires measurements of individual particle contacts and vector forces. Recent developments in photoelastic techniques [30, 31, 32], combined with high-resolution photography, now allow for force and contact resolution suitable to test the SFF relation using photoelastic granular assemblies.
@@
In this study, we use photoelastic disks  [30, 31] arranged in a single layer within an annular shear cell (Fig. 1(a)) to map emergent particle scale anisotropic structures to bulk friction and the stress tensor using the SFF relation and further apply it to capturing dynamic behaviour. In doing so, we address the open question of the influence of particle scale anisotropy during granular dynamics and show the first laboratory demonstration of the SFF relation. The particles begin in an unstrained state with no observable forces and the inner boundary is rotated to shear the assembly. During these experiments, we record particle locations, contacts, and normal and frictional forces vectors for each particle and calculate the stress tensor and anisotropies. In the companion article we use the SFF relation to study the role of loading condition on the evolution of bulk properties [33]
@@
We performed experiments in a constant-area quasi-two-dimensional annular shear cell containing $N_{p}$ = 1810 bidisperse (Vishay PhotoStress PSM-4) circular disks with an equal fraction of small particles with radii $R_{s}$ = 4.5 mm and large particles with radii $R_{l}$ = 5.5 mm. The particles are dusted with baking powder to reduce basal friction between the particles and the supporting sheet, and the interparticle friction coefficient is estimated to be 0.3. The inner and outer boundaries of the annulus have particle-sized cavities to promote gripping between the boundaries and the particles. The inner boundary of the annulus is rotated at a constant rate of 1 rotation every 2.5 hours and the packing undergoes simple shear. Further details of the annulus can be found in previous works [34, 35]. Using the annular strain rate $\dot{\gamma}=\frac{1}{2}\left(\frac{\partial v}{\partial r}-\frac{v}{r}\right)$, and integrating over time with appropriate boundary conditions, we calculate the strain $\epsilon$. The packing is illuminated with both red unpolarized light (particle-detection) and green circularly polarized light (darkfield polariscope). Images are taken at a frame rate of 5 seconds per frame with a Nikon D850 camera. We extract particle position, contact and vector force data from the images of the particles as outlined in previous works  [30, 31] by using the open source code, PhotoElastic Grain Solver [32] that we modified to detect particle-particle contacts with force magnitudes with a resolution of $5\times 10^{-5}$ N. Further details of the analysis can be found in the companion article [33].
@@
From the particle scale information, we calculate the stress tensor in two dimensions using the micro-structural definition [36, 37]. We compute the two dimensional stress tensor via summation of the outer product of the force $\vec{f}$ and contact position vector $\vec{r}$ [37, 29, 23]:
$\sigma_{ij}=\frac{1}{S}\sum_{a\neq b}\vec{r^{ab}_{i}}\vec{f^{ab}_{j}},$ (1)
@@
where in our two dimensional system, $S$ is the area of the annulus. Here we follow classical notation, where the contact position vector points from the center of the grain to the contact point, and the force vector points from the contact point in the direction of the contact force. Figure 1(b) shows a schematic of the vectors for two sample particles. From this formulation of the stress tensor, we are able to calculate the bulk friction coefficient,
@@
where the numerator is the shear component of the stress tensor, and the denominator is the mean pressure.
To consider the anisotropies in a granular material, the SFF relation relies on decomposing the stress tensor into angular distributions according to the unit contact vector $\hat{n}$. In 2D, the unit contact vector is described by a single angle $\theta$, and assuming the thermodynamic limit, the stress tensor can be written as:
[…]
where $N_{c}$ is the total number of contacts in $S$, $\vec{\ell}$ is the branch vector (the vector joining the centers of two contacting particles), and $E^{c}({\theta}$) is the probability distribution function (pdf) of a contact pointing along the direction $\theta$.
Because our system is composed of round, nearly monodisperse particles, the branch vector can be approximated as $\vec{\ell}=d\hat{n}$, with $d$ the mean grain diameter, and therefore there is no preferred orientation to the branch vector [25]. Because, in general, the vector forces and branches are statistically independent, we approximate $\langle\ell_{i}f_{j}\rangle(\theta)\sim dn_{i}(\theta)\langle f_{j}\rangle(\theta)$. The angular distribution of the mean vector force $\langle f_{j}\rangle(\theta)$ can be decomposed according the mean angular distribution of normal $\langle f_{n}\rangle(\theta)$ and tangential $\langle f_{t}\rangle(\theta)$ forces as $\langle f_{j}\rangle(\theta)=\langle f_{n}\rangle(\theta)n_{j}+\langle f_{n}\rangle(\theta)t_{j}$.
Figure 2: Normalized angular distributions of the (a) contacts $E^{c}(\theta)$, (b) normal forces $\langle f_{n}\rangle(\theta)$, and (c) tangential forces $\langle f_{t}\rangle(\theta)$ at increasing levels of strain $\epsilon$ (dark purple to yellow). The contacts are normalized by the average number of contacts $E_{0}$, and the forces are normalized by the mean normal force $f_{0}$ at each given strain. Sample angular histograms of each anisotropic metric are shown in the respective insets, which have been averaged over the strain range shown in the main panel. $\theta=0$ corresponds to the radial direction, and $\theta=\pi/2$ corresponds to the azimuthal direction. Fits to the respective Fourier expansions are indicated by the solid, sinusoidal line (Eqs. 4(a)-(c)). The normalized contacts and normalized normal forces show a decrease in anisotropy with increasing strain.
In general, these distributions are approximated by a first-order Fourier expansion such that [23, 38, 29]:
@@
$\langle f_{n}\rangle(\theta)$ $=f_{0}\left[1+a_{n}\cos 2\left(\theta-\theta_{a}\right)\right],$ (4b)
$\langle f_{t}\rangle(\theta)$ $=f_{0}a_{t}\sin 2\left(\theta-\theta_{a}\right),$ (4c)
@@
where $\theta_{a}$ is the angle corresponding to the principle stress direction, $f_{0}$ is the average magnitude of the normal force, and $a,a_{n},a_{t}$ are the amplitudes of the leading order Fourier modes for the contact, normal force and tangential force distributions, respectively.
@@
Examples of the three distributions presented in Eqs 4(a)-(c) are shown in Fig. 2 as polar plots averaged over the entire deformation (inset) and using a colour scale at varying strain indicated from purple (dark) to yellow (light) in increments of 0.04. We plot the change in these anisotropic measures as a function of relative angle between the radial direction and branch vector, where $\theta=0$ indicates alignment with the radial direction and $\theta=\pi/2$ indicates alignment with the azimuthal direction. We expect the primary direction of the normal forces and contacts to align with the $\theta\sim\pi/4$ direction as the load is distributed through the shear band. This is true for the normal forces, indicating that the majority of the load is transmitted in the shear direction. The contact distribution follows a similar trend, however there is an additional secondary feature that appears as a bump aligned in the azimuthal direction $\theta=\pi/2$. Two possible reasons for the increased number of contacts near the azimuthal direction is either due to an artifact during contact detection, or an indication of a higher order Fourier mode, possibly present due to the constant area boundary condition. Regardless, this feature is small relative to the main signal, and we use only one Fourier mode to describe the data. The tangential distributions show only a weak signal above the noise, indicating that there is no strong preference for average tangential forces to be either clockwise (purple) or counterclockwise (pink) from the contact normal.
Following [23, 29], we expand the stress tensor (Eq. 3) with the Fourier approximations (Eqs. 4 (a)-(c)) to compare the particle scale anisotropies to bulk metrics. In addition, we substitute the number of contacts per unit area as $N_{c}/S=2z\nu/(\pi d^{2})$, where $z=2N_{c}/N_{p}$ is the coordination number, $\nu$ the packing fraction and $d$ the mean grain diameter. The SFF relation states that the stress tensor can be re-written in terms of metrics related to the topology of the contact and forces network as
@@
$\sigma_{ij}={\frac{z\nu}{\pi d}f_{0}}\,\Biggl{[}\left(1+\frac{aa_{n}}{2}\right)\delta_{ij}\\ +\frac{1}{2}\left(a+a_{n}+a_{t}\right)\begin{pmatrix}\cos 2\theta_{a}&\sin{2}\theta_{a}\\ \sin 2\theta_{a}&-\cos{2}\theta_{a}\end{pmatrix}\Biggr{]},$ (5)
Then, inserting Eq. 5 into Eq. 2, and neglecting cross product between the anisotropies, $\mu$ results in a simple sum rule:
@@
which directly relates the bulk friction coefficient to the anisotropy amplitude, indicating that the resistance to deformation in a granular material simply depends on the amount of anisotropy that is present in the granular assembly.
[b]
Figure 3: Panels (a)-(c) display the evolution of the packing fraction $\nu$, coordination number $z$ with (red) rattlers and without (cyan) rattlers (i.e., particles with less than 2 contacts) and the mean normal force $f_{0}$, respectively, as a function of strain $\epsilon$. Panels (d)-(f) display the leading term Fourier coefficients representing the angular distribution of contacts ($a$), average normal forces ($a_{n}$), and tangential forces ($a_{t}$), respectively, within a strained granular packing under shear. These coefficients quantify the degree of anisotropy observed in each metric within the packing.
To consider the microstructure of the packing, we measure the first order metrics of connectivity. Fig.3(a)-(c) shows the evolution of the packing fraction $\nu$, coordination number $z$ (with or without considering rattlers) and mean normal force $f_{0}$ as a function of strain. The solid fraction can be considered as constant throughout the strain, while $z$ with (red) or without (cyan) rattlers first increases with strain before saturating around a value of $2$, $3$, respectively. Finally, $f_{0}$ follows the same trend, first increasing before tending to a constant value in the steady state (also referred as “critical state” in soil mechanics).
@@
To estimate the fabric and force networks anisotropies, we fit the angular distribution of contacts, normal forces and tangential forces (Fig. 2) to the predicted Fourier decomposition in Eqs. 4(a)-(c). We take the principle stress direction $\theta_{a}$ and the amplitudes $a$, $a_{n}$ and $a_{t}$ to be free fit parameters, where the anisotropy corresponds to the amplitudes. The anisotropies are plotted as a function of strain in Fig. 3(d)-(f); note that during the initial strain there is a strong decrease in value for all three metrics, indicating a decrease in overall anisotropy. At larger strains the three amplitudes tend toward a constant value, meaning that a steady state is reached. In the steady state the relative magnitude of $a$ and $a_{n}$ are comparable, which is consistent with simulated systems of disks [25, 23], while $a_{t}$ is an order of magnitude smaller. This indicates that the tangential forces, and consequently the inter-particle friction, has only a small effect in our system on the bulk friction, due to $a_{t}$ being smaller than those reported in simulations [26, 25].
Figure 4: Panel (a) illustrates the bulk friction coefficient of the sheared granular packing as a function of shear, with the summed anisotropies overlaid in black and the standard error in the fit represented as grey filled bars spanning the best fit. The fabric anisotropy $a$ is plotted with a black dashed line for comparison. The blue panel highlights the strain $\epsilon$ depicted in fig.3 (a)-(f). Panel (b) presents the four components of the strain tensor, $\sigma$, as measured by Eq. 1, with principal and secondary stress axes oriented radially and azimuthally, respectively. The solid lines represent the stress tensor calculated from the anisotropies using Eq. 5.
@@
Figure 4(a) shows the evolution of $\mu$, measured directly from the components of the granular stress tensor (Eq.2) and as predicted by Eq.6. The prediction from the SFF relation is excellent throughout the deformation and matches the measured value even through large rearrangements and stick slip events which are indicated by large spikes in $\mu$. This result shows for the first time an experimental test of the SFF relationship. For an annular shear cell, we expect $\mu$ to initially decrease as it undergoes strain softening before reaching a steady state [39]. Based on the SFF relation, this indicates that the fabric and forces are initially more anisotropic before becoming less anisotropic with increased strain. Notably, $\mu$ is not solely determined by the number of forces (see supplemental video [40]), as there are strong fluctuations in the number of contact forces for a given value of $\mu$.
@@
To explore the applicability of the SFF relation to systems without contact force data, in Fig. 4(a) we show the effect of the fabric anisotropy alone by plotting only $a$ as a function of strain (dashed black line). Overall, $a$ follows the overall pattern of $\mu$, including the initial transient and stick-slip events, but underestimates bulk friction by roughly 30%. Because $a$ and $a_{n}$ are typically of the same magnitude, while the tangential component is small, using the fabric anisotropy alone provides a good estimate of bulk friction at least for circular particles with low polydispersity [25, 41].
As well, we plot the stress tensor components, as a function of strain, in Fig 4(b). The radial and azimuthal stress increase during strain, and then fluctuate as the particles undergo rearrangements during stick slip events. The off-axis components of the stress tensor remain small. Overlaid on the calculated stress tensor components are the predictions for the stress tensor based on Eq. 5. Again, these agree well with the measured values. To gain a physical understanding of the influence of the microscale on the stress tensor, we replace the sum of anisotropies by $\mu$ in Eq.5, and neglect the cross product $a\,a_{n}$. The rapid increase in $\sigma_{RR}$ and $\sigma_{\phi\phi}$, mainly results from the increases in $z$ and $f_{0}$. Moreover, in the steady state $z\nu f_{0}(1+\mu)/(\pi d)$ provide a fairly good approximation of both $\sigma_{RR}$ and $\sigma_{\phi\phi}$.
@@
In this study, we have shown an experimental validation of the SFF relation using photoelastic particles in an annular shear cell. Success of the SFF relation indicates that the resistance to deformation of a granular material arises due to anisotropies in fabric and forces. Importantly, the SFF relation remains valid even as transient stress builds up and releases through stick slip events, far beyond the static loading for which the framework was initially developed. We note that this framework does not address the role of mesoscale structure and therefore does not encompass prediction of rigidity [5]; however, the SFF relation is a promising pathway to address the open challenge of connecting fluid-like granular theories [14, 13, 15] to quasi-static granular behaviour. Finally, the framework may additionally provide a way to link fabric anisotropy to bulk friction in experimental systems for which vector contact forces cannot be measured, as illustrated by the success of the contact anisotropy $a$ alone. We note that in the absence of force information, the geometric anisotropy provides a good estimate of bulk friction,at least when slightly polydisperse discs are considered [25, 26], and could be easily applied to other deformable assemblies [42], including cellular tissues, foams, and dense colloid suspensions.